Learning Goal:

To calculate average and relative reaction rates.

You can measure the rate of a reaction, just like you can measure the speed a jogger runs. While a jogger would be reported to run a specific number of miles in an hour, miles/hour, a reaction is reported to form product or consume reagent in molar concentration per second, M/s.

Reaction rate can be defined either as the increase in the concentration of a product per unit time or as the decrease of the concentration of a reactant per unit time. By definition, reaction rate is a positive quantity.

In the reaction X→2Y, for example, Y is being produced twice as fast as X is consumed and thus

rate of X=12(rate of Y)

Each rate can be expressed as the change in concentration over the change in time, Δt:

−Δ[X]Δt=12(Δ[Y]Δt)

Note: Although the rate of changes in specific concentrations can be assigned positive or negative values depending on if a product or reactant, respectively, reaction rates always have positive values.

Part A

Consider the reaction

2H3PO4→P2O5+3H2O

Using the information in the following table, calculate the average rate of formation of P2O5 between 10.0 and 40.0 s.

Time (s) 0 10.0 20.0 30.0 40.0 50.0
[P2O5] (M) 0 3.00×10−3 6.00×10−3 7.80×10−3 9.00×10−3 9.60×10−3

Express your answer with the appropriate units.

Rate of formation of P2O5 =

Part B

Complete previous part(s)

Part C

Consider the reaction

5Br−(aq)+BrO−3(aq)+6H+(aq)→3Br2(aq)+3H2O(l)

The average rate of consumption of Br− is 1.56×10−4 M/s over the first two minutes. What is the average rate of formation of Br2 during the same time interval?

Express your answer with the appropriate units.

Part D

Complete previous part(s)
Answer #1

Page: Given reaction is 2H₂ Poy Post 3 H2O buse Rate of fooation of P205 between los and uos = 9x10-3 3x10-3 m 305 = 6X longDU Page: --3 (1-1.56x154 m) = f (852) olevas (1:56 x 100km) 29.36x105m rate of formation of Boy is 9.36x10-5